OpenStudy (anonymous):

Please Hellpp!! Suppose you want to join a video store. Big Video offers a special discount card that costs $9.99 for one year. With the discount card, each video rental costs $2.49. A discount card from Main Street Video costs $20.49 for one year. With the Main Street Video discount card, each video rental costs $1.79. After how many video rentals is the cost the same?

OpenStudy (anonymous):

Suppose you are thinking about buying one of two cars. Car A will cost $17,655. You can expect to pay an average of $1230 per year for fuel, maintenance, and repairs. Car B will cost about $15,900. Fuel, maintenance, and repairs for it will average about $1425 per year. After how many years are the total costs for the cars the same?

OpenStudy (mr.math):

The cost of x video rentals in Big Vid is \(c_1(x)=9.99+2.49x\) and in Main Street Vid is \(c_2(x)=20.49+1.79x\). Now to get when they are equal, solve the linear equation: \[9.99+2.49x=20.49+1.79x\]

OpenStudy (mr.math):

Which I'm sure you can do by yourself.